# A Methodology to Study the Impact of Spiking Neural Network Parameters considering Event-Based Automotive Data

### v2

@@ I Introduction
In recent years, the interest in the autonomous driving (AD) systems has increased significantly as these systems potentially improve the efficiency of human mobility and transportation [1][2]. The development of AD systems in real life necessitates the capabilities of solving computer vision tasks such as image classification and object detection/segmentation from images/videos [3]. Currently, the state-of-the-art accuracy for computer vision tasks is achieved through advanced neural network algorithms, such as Deep Neural Networks (DNNs) and Spiking Neural Networks (SNNs). Therefore, the employment of neural network algorithms for AD systems is prevalent [4]. Besides the above-discussed functionalities, AD systems also require low energy consumption to preserve their battery lifespan, and real-time output to provide fast decision. These requirements can potentially be fulfilled by SNNs, since SNNs offer low power/energy computation due to sparse event-based operations [5, 6, 7], high accuracy due to effective learning mechanism [1][2][8][9], and low latency due to efficient neural/spike coding and operational timesteps [10, 11, 12, 13]. Furthermore, to maximize the benefits of sparse operations in SNNs, the workload should be presented in the form of event-based data, because this data format is suitable with the data representation in SNNs (i.e., spikes). Therefore, to ensure the practicality of the SNN models for AD systems, the SNN developments should also consider event-based automotive data, such as the NCARS dataset [14][15]. Motivated by the potentials of SNN-based AD systems, our targeted research problems are the following.
  • How can we investigate, analyze, and understand the impact of different SNN parameters on the final accuracy?
@@ I-B Our Novel Contributions
To address the above-discussed problems and challenges, we propose a novel methodology to systematically analyze the impact of SNN parameters on the learning quality, and then leverage this analysis for devising effective SNN architectures for AD systems. Its key steps are presented in the following (see an overview in Fig. 2).
@@
  • Improving SNN learning quality based on our parameter enhancements (Section III-C): This step leverages the knowledge from previous analysis by tuning the parameter settings to improve the SNN accuracy and/or reduce the training time, hence leading to better SNNs for AD systems.
@@ II-C SNN Architecture
The SNN architecture considered in this work is based on the studies in [1], and supports the STBP-based learning mechanism. The architectural details are provided in the Table I. It is composed of two convolutional layers with 32 output channels, $3\times 3$ kernel size, padding of 1, and stride of 1, interleaved by average pooling layers, and followed by two fully-connected layers (a.k.a. dense layers).
@@ III-B Exploring the Impact of Different SNN Parameter Values
It aims at exploring different settings (values) of the selected SNN parameters, and analyzes their impact on the accuracy. Here, each parameter has its own range of exploration settings due to the nature of their functionalities. To decide the range of values for each parameter, we first consider the parameter settings in the state-of-the-art works as the default settings. For instance, the default setting for the SNN architecture with STBP rule [1] considers the following set of values: $B=40$, $lr=1e^{-3}$, $V_{th}=0.4$, and $w_{decay}=0$. The experimental results for employing the default setting are presented in Fig. 6. The results show that, in general, the accuracy curve consists of two regions: (1) transition region, where the accuracy is still not stable as the network is still learning new information over training time, and (2) stable region, where the accuracy does not change much as the network does not necessarily learn new information. This curve indicates how fast the training phase reaches stable accuracy. We define the stable region as the region that has accuracy with a standard deviation of less than 0.5% accuracy over the last 10 training epochs, since the typical range of acceptable accuracy variation in the neural network community is within 1% accuracy [22][23].
Figure 6: The experimental results of the accuracy under the default setting: $B=40$, $lr=1e^{-3}$, $V_{th}=0.4$, and $w_{decay}=0$. We observe that the accuracy curve typically consists of two regions, i.e., transition and stable.
@@
  • $lr\in\{5e^{-4},1e^{-4},1e^{-3},5e^{-3},7.5e^{-3},1e^{-2},5e^{-2}\}$.
@@
Based on these settings, we perform experiments that tune each investigated parameter with the desired value at one time, while keeping other parameters set with the default values. For instance, if we investigate the impact of $B=10$, then we perform experiments with the following setting: $B=10$, $lr=1e^{-3}$, $V_{th}=0.4$. Note, the experimental results and related discussion for the exploration will be provided in Section V-A until Section V-D.
III-C Parameter Enhancements for Improving the SNN Learning Quality
@@ IV Evaluation Methodology
Fig. 7 illustrates the experimental setup for evaluating our methodology. Here, we employ a Python-based implementation that runs on an Nvidia RTX 6000 Ada GPU machine, and the generated outputs are the accuracy (i.e., training and test) and the log of experiments (e.g., number of epoch and loss scores). The Python-based implementation is built using the numpy, torch, and torchvision library packages. We employ the SNN architecture described in Table I with the STBP learning rule, the NCARS dataset as the workload, and 200 epochs for the training phase. We consider the state-of-the-art work, i.e., CarSNN [1] with the following default setting: $B=40$, $lr=1e^{-3}$, $V_{th}=0.4$, and $w_{decay}=0$, as the reference comparison partner.
@@ V-B Impact of the Learning Rate
The experimental results for the impact of learning rate $lr$ are presented in Fig. 9. The results show that, in general, most of the learning rates have fluctuating accuracy during the early training phase (i.e., $\leq$ 60 epoch) since the network still starts learning new information; see label  3. In this early training phase, we observe that learning rates $7.5e^{-3}$ and $1e^{-2}$ can quickly reach accuracy without notable fluctuations, while smaller and bigger learning rates face significant fluctuations. The reason is that, smaller and bigger learning rates struggle to find local minima, while the learning rates of $7.5e^{-3}$ and $1e^{-2}$ are in the right range of values to generalize well. These trends stay consistent until later training phase, as shown by label  4, indicating that the learning rates of $7.5e^{-3}$ and $1e^{-2}$ are effective in training the SNN to high accuracy values compared to other learning rates. To decide which learning rate value to select as the adjustment setting, we compare the accuracy obtained by learning rates $7.5e^{-3}$ and $1e^{-2}$. Overall, Fig. 9 shows that a learning rate of $7.5e^{-3}$ achieves comparable accuracy to the learning rate of $1e^{-2}$ over the training phase, and in the range of 20-40 training epochs, the learning rate of $37.5e-$ 3 has slightly better results, meaning that it potentially enables a faster learning process. Therefore, in this work, we select $lr=7.5e^{-3}$ as the effective value for SNN enhancements. Later, we will also show in Section V-F that employing $lr=1e^{-2}$ can also achieve good accuracy, demonstrating the effectiveness of our methodology.
Figure 9: Experimental results for the accuracy across different learning rates, i.e., $lr\in\{5e^{-4},1e^{-4},1e^{-3},5e^{-3},7.5e^{-3},1e^{-2},5e^{-2}\}$. Here, we observe that learning rates of $7.5e^{-3}$ and $1e^{-2}$ can achieve better accuracy than other learning rates.
@@ V-E Impact of the SNN Parameter Enhancements
Based on previous analysis from Section V-A until Section V-D, we propose a set of parameter values for enhancing SNNs for an event-based automotive dataset: $B=20$, $lr=7.5e^{-3}$, $V_{th}=0.5$, and $w_{decay}=0$. The experimental results for our parameter enhancements are presented in Fig. 12. The results show that, in general, our parameter enhancements improve the SNN accuracy from the state-of-the-art over the training epoch, demonstrating the effectiveness of our parameter settings. Our enhanced SNN model has a better accuracy range over the training phase (71.28% - 86.64% accuracy) than the state-of-the-art which has a range of 57.06% - 85.84% accuracy. In the early training phase, the state-of-the-art work faces significant accuracy fluctuations as its parameter values are crafted to achieve stable accuracy after running hundreds of training epochs; see label  8. Meanwhile, our parameter enhancements can achieve a relatively smooth learning curve (i.e., accuracy) in the transition region, since we craft its parameter values with the ones that iteratively find the local minima over the training phase; see label  8. These trends stay consistent until the later training phase, since our parameter enhancements achieve notable accuracy improvements as compared to the state-of-the-art work; see label  9. Furthermore, our proposed parameter enhancements can reach a stable region faster than the state-of-the-art work (i.e., by 1.9x speed up). Here, the state-of-the-art work achieves 85.24% accuracy after performing 125 training epochs, while our enhanced SNN model achieves 85.99% accuracy after performing 66 training epochs; see label  10. The reason for all these improvements is that we employ parameter enhancements that combine benefits from the selected values, thus maximizing their advantages together for improving SNN accuracy and training time.
Figure 12: Comparison of accuracy obtained by the state-of-the-art work and our proposed methodology, with the following setting: $B=20$, $lr=7.5e^{-3}$, $V_{th}=0.5$, and $w_{decay}=0$. The proposed parameter enhancements lead to higher accuracy than the state-of-the-art work, with a better learning curve over the training phase.
@@ V-F Further Discussion
  • Setting-1: $B=20$, $lr=7.5e^{-3}$, $V_{th}=0.5$, $w_{decay}=0$.
  • Setting-2: $B=20$, $lr=1e^{-2}$, $V_{th}=0.5$, $w_{decay}=0$.
Setting-1 is simply the same as our proposed one discussed in the previous subsection (i.e., Section V-E), while Setting-2 is similar to Setting-1 but with a different learning rate. We select a learning rate of $1e^{-2}$ for Setting-2 because from the observations in Section V-B, this learning rate can achieve comparable accuracy to our selected learning rate value (i.e., $7.5e^{-3}$). The experimental results for this case study are presented in Fig. 13. The results show that, in general, Setting-2 leads the enhanced SNN model to achieve a better accuracy range over the training phase (i.e., 75.48% - 86.46% accuracy) than the state-of-the-art work as indicated by label  11 and  12, demonstrating the effectiveness of its parameter setting. Furthermore, Setting-2 can also reach a stable region faster than the state-of-the-art work with 85.17% accuracy, similar to that of Setting-1. The reason for all these improvements is that the SNN parameter values in Setting-2 are also obtained through our methodology, hence combining benefits from the selected parameter values for improving SNN models with event-based autonomous data.
Figure 13: Comparison of accuracy obtained by the state-of-the-art work and our proposed methodology considering different settings, i.e., (1) Setting-1: $B=20$, $lr=7.5e^{-3}$, $V_{th}=0.5$, $w_{decay}=0$, and (2) Setting-2: $B=20$, $lr=1e^{-2}$, $V_{th}=0.5$, $w_{decay}=0$.
@@ Acknowledgment
This work has been supported in parts by the NYUAD Center for Cyber Security (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104, and Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.

### v3

@@ I Introduction
In recent years, the interest in the autonomous driving (AD) systems has increased significantly as these systems potentially improve the efficiency of human mobility and transportation [1][2]. The development of AD systems in real life necessitates the capabilities of solving computer vision tasks such as image classification and object detection/segmentation from images/videos. Currently, the state-of-the-art accuracy for computer vision tasks is achieved through advanced neural network algorithms, such as Deep Neural Networks (DNNs) and Spiking Neural Networks (SNNs). Therefore, the employment of neural network algorithms for AD systems is prevalent [3]. Besides the above-discussed functionalities, AD systems also require low energy consumption to preserve their battery lifespan, and real-time output to provide fast decision. These requirements can potentially be fulfilled by SNNs, since SNNs offer low power/energy computation due to sparse event-based operations [4, 5, 6], high accuracy due to effective learning mechanism [1][2][7][8], and low latency due to efficient neural/spike coding and operational timesteps [9, 10, 11, 12]. Furthermore, to maximize the benefits of sparse operations in SNNs, the workload should be presented in the form of event-based data, because this data format is suitable with the data representation in SNNs (i.e., spikes). Therefore, to ensure the practicality of the SNN models for AD systems, the SNN developments should also consider event-based automotive data, such as the NCARS dataset [13]. Motivated by the potentials of SNN-based AD systems, our targeted research problems are the following.
  • How can we investigate, analyze, and understand the impact of different SNN parameters on final accuracy?
@@ I-B Our Novel Contributions
To address the above-discussed problems and challenges, we propose a novel methodology to systematically analyze the impact of SNN parameters on the learning quality, and then leverage this analysis for devising effective SNN architectures for AD systems. Its key steps are presented in the following (see overview in Fig. 2).
@@
  • Improving SNN learning quality based on our parameter enhancements (Section III-C): This step leverages the knowledge from previous analysis by tuning the parameter settings to improve the SNN accuracy and/or reduce the training time, thus leading to better SNNs for AD systems.
@@ II-C SNN Architecture
The SNN architecture considered in this work is based on the studies in [1], and supports the STBP-based learning mechanism. The architectural details are provided in the Table I. It is composed of two convolutional layers with 32 output channels, $3\times 3$ kernel size, padding of 1, and stride of 1, interleaved by average pooling layers, and followed by two fully-connected layers (dense layers).
@@ III-B Exploring the Impact of Different SNN Parameter Values
It aims at exploring different settings (values) of the selected SNN parameters, and analyzes their impact on the accuracy. Here, each parameter has its own range of exploration settings due to the nature of their functionalities. To decide the range of values for each parameter, we first consider the parameter settings in the state-of-the-art works as the default settings. For instance, the default setting for the SNN architecture with STBP rule [1] considers the following set of values: $B=40$, $lr=1e\text{-}3$, $V_{th}=0.4$, and $w_{decay}=0$. The experimental results for employing the default setting are presented in Fig. 6. The results show that, in general, the accuracy curve consists of two regions: (1) transition region, where the accuracy is still not stable as the network is still learning new information over training time, and (2) stable region, where the accuracy does not change much as the network does not necessarily learn new information. This curve indicates how fast the training phase reaches stable accuracy. We define the stable region as the region that has accuracy with a standard deviation of less than 0.5% accuracy over the last 10 training epochs, since the typical range of acceptable accuracy variation in the neural network community is within 1% accuracy [20][21].
Figure 6: The experimental results of the accuracy under the default setting: $B=40$, $lr=1e\text{-}3$, $V_{th}=0.4$, and $w_{decay}=0$. We observe that the accuracy curve typically consists of two regions, i.e., transition and stable.
@@
  • $lr\in\{5e\text{-}4,1e\text{-}4,1e\text{-}3,5e\text{-}3,7.5e\text{-}3,1e\text{-}2,5e\text{-}2\}$.
@@
Based on these settings, we perform experiments that tune each investigated parameter with the desired value at one time, while keeping other parameters set with the default values. For instance, if we investigate the impact of $B=10$, then we perform experiments with the following setting: $B=10$, $lr=1e\text{-}3$, $V_{th}=0.4$. Note, the experimental results and related discussion for the exploration will be provided in Section V-A until Section V-D.
III-C Parameter Enhancements for Improving SNN Learning
@@ IV Evaluation Methodology
Fig. 7 illustrates the experimental setup for evaluating our methodology. Here, we employ a Python-based implementation that runs on an Nvidia RTX 6000 Ada GPU machine, and the generated outputs are the accuracy (i.e., training and test) and the log of experiments (e.g., number of epoch and loss scores). The Python-based implementation is built using the numpy, torch, and torchvision library packages. We employ the SNN architecture described in Table I with the STBP learning rule, the NCARS dataset as the workload, and 200 epochs for the training phase. We consider the state-of-the-art work, i.e., CarSNN [1] with the following default setting: $B=40$, $lr=1e\text{-}3$, $V_{th}=0.4$, and $w_{decay}=0$, as the reference comparison partner.
@@ V-B Impact of the Learning Rate
The experimental results for the impact of learning rate $lr$ are presented in Fig. 9. The results show that, in general, most of the learning rates have fluctuating accuracy during the early training phase (i.e., $\leq$ 60 epoch) since the network still starts learning new information; see label  3. In this early training phase, we observe that learning rates $-37.5e\text{-}$ 3 and $-21e\text{-}$ 2 can quickly reach accuracy without notable fluctuations, while smaller and bigger learning rates face significant fluctuations. The reason is that, smaller and bigger learning rates struggle to find local minima, while the learning rates of $-37.5e\text{-}$ 3 and $-21e\text{-}$ 2 are in the right range of values to generalize well. These trends stay consistent until later training phase, as shown by label  4, indicating that the learning rates of $-37.5e\text{-}$ 3 and $-21e\text{-}$ 2 are effective in training the SNN to high accuracy values compared to other learning rates. To decide which learning rate value to select as the adjustment setting, we compare the accuracy obtained by learning rates $-37.5e\text{-}$ 3 and $-21e\text{-}$ 2. Overall, Fig. 9 shows that a learning rate of $-37.5e\text{-}$ 3 achieves comparable accuracy to the learning rate of $-21e\text{-}$ 2 over the training phase, and in the range of 20-40 training epochs, the learning rate of $-37.5e\text{-}$ 3 has slightly better results, meaning that it potentially enables a faster learning process. Therefore, in this work, we select $lr=7.5e\text{-}3$ as the effective value for SNN enhancements. Later, we will also show in Section V-F that employing $lr=1e\text{-}2$ can also achieve good accuracy, demonstrating the effectiveness of our methodology.
Figure 9: Experimental results for the accuracy across different learning rates, i.e., $lr\in\{5e\text{-}4,1e\text{-}4,1e\text{-}3,5e\text{-}3,7.5e\text{-}3,1e\text{-}2,5e\text{-}2\}$. Here, we observe that learning rates of $-37.5e\text{-}$ 3 and $-21e\text{-}$ 2 can achieve better accuracy than other learning rates.
@@ V-E Impact of the SNN Parameter Enhancements
Figure 12: Comparison of accuracy obtained by the state-of-the-art work and our proposed methodology, with the following setting: $B=20$, $lr=7.5e\text{-}3$, $V_{th}=0.5$, and $w_{decay}=0$. The proposed parameter enhancements lead to higher accuracy than the state-of-the-art work, with a better learning curve over the training phase.
Based on previous analysis from Section V-A until Section V-D, we propose a set of parameter values for enhancing SNNs for an event-based automotive dataset: $B=20$, $lr=7.5e\text{-}3$, $V_{th}=0.5$, and $w_{decay}=0$. The experimental results for our parameter enhancements are presented in Fig. 12. The results show that, in general, our parameter enhancements improve the SNN accuracy from the state-of-the-art over the training epoch, demonstrating the effectiveness of our parameter settings. Our enhanced SNN model has a better accuracy range over the training phase (71.28% - 86.64% accuracy) than the state-of-the-art which has a range of 57.06% - 85.84% accuracy. In the early training phase, the state-of-the-art work faces significant accuracy fluctuations as its parameter values are crafted to achieve stable accuracy after running hundreds of training epochs; see label  8. Meanwhile, our parameter enhancements can achieve a relatively smooth learning curve (i.e., accuracy) in the transition region, since we craft its parameter values with the ones that iteratively find the local minima over the training phase; see label  8. These trends stay consistent until the later training phase, since our parameter enhancements achieve notable accuracy improvements as compared to the state-of-the-art work; see label  9. Furthermore, our proposed parameter enhancements can reach a stable region faster than the state-of-the-art work (i.e., by 1.9x speed up). Here, the state-of-the-art work achieves 85.24% accuracy after performing 125 training epochs, while our enhanced SNN model achieves 85.99% accuracy after performing 66 training epochs; see label  10. The reason for all these improvements is that we employ parameter enhancements that combine benefits from the selected values, thus maximizing their advantages together for improving SNN accuracy and training time.
@@ V-F Further Discussion
  • Setting-1: $B=20$, $lr=7.5e\text{-}3$, $V_{th}=0.5$, $w_{decay}=0$.
  • Setting-2: $B=20$, $lr=1e\text{-}2$, $V_{th}=0.5$, $w_{decay}=0$.
Setting-1 is simply the same as our proposed one discussed in the previous subsection (i.e., Section V-E), while Setting-2 is similar to Setting-1 but with a different learning rate. We select a learning rate of $-21e\text{-}$ 2 for Setting-2 because from the observations in Section V-B, this learning rate can achieve comparable accuracy to our selected learning rate value (i.e., $-37.5e\text{-}$ 3). The experimental results for this case study are presented in Fig. 13. The results show that, in general, Setting-2 leads the enhanced SNN model to achieve a better accuracy range over the training phase (i.e., 75.48% - 86.46% accuracy) than the state-of-the-art work as indicated by label  11 and  12, demonstrating the effectiveness of its parameter setting. Furthermore, Setting-2 can also reach a stable region faster than the state-of-the-art work with 85.17% accuracy, similar to that of Setting-1. The reason for all these improvements is that, the SNN parameter values in Setting-2 are also obtained through our methodology, hence combining benefits from the selected parameter values for improving SNN models with event-based autonomous data.
Figure 13: Comparison of accuracy obtained by the state-of-the-art work and our proposed methodology considering different settings, i.e., (1) Setting-1: $B=20$, $lr=7.5e\text{-}3$, $V_{th}=0.5$, $w_{decay}=0$, and (2) Setting-2: $B=20$, $lr=1e\text{-}2$, $V_{th}=0.5$, $w_{decay}=0$.
@@ Acknowledgment
This work was partially supported by the NYUAD Center for Interacting Urban Networks (CITIES), funded by Tamkeen under the NYUAD Research Institute Award CG001.